Find the equation of the line that passes through the point (3,-1) and is parallel to the line y=-x+4
Answer by ikleyn(52781)

.

The line parallel to y = x+4  has the form  y = x + c,  where "c" is  some constant.


We find the value of "c" by substituting the coordinates of the given point (3,-1) into the equation

    -1 = 3 + c,


which gives  c = -1 - 3 = -4.


So, the equation of the requested line is  

    y = x -4.      ANSWER

Solved and completed.

-------------------

There is a group of closely related lessons in this site

    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point
    - Advanced problems on finding equations for straight lines

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.